OpenStudy (anonymous):

I have just 2 more problems I need help with. Expand the following logarithmic expression: log (3x^2/12) The answer I came up with is log 3+2 log x-log 12, is this correct? Second question: Simplify the following expression: log_a b^2. I come up with 2log_ab, is this right?

OpenStudy (anonymous):

to make more sense I will write the actual equation for the second question. \[\log _{a}b ^{2}\]

Parth (parthkohli):

The second is right. I don't get your solution for the first. =\

OpenStudy (jiteshmeghwal9):

\(log_ab-log_ac=log_a\frac{b}{c}\)

OpenStudy (jiteshmeghwal9):

\[\log(3x^2/12)=\log(3x^2)-\log(12)\]

Parth (parthkohli):

Yes ^

OpenStudy (anonymous):

here is the equation for number one \[\log(\frac{ 3x ^{2} }{ 12 })\]

OpenStudy (jiteshmeghwal9):

\[\log_ab^c=clog_ab\]so ur second answer is right :)

OpenStudy (jiteshmeghwal9):

use the property\[\log_ab-\log_ac=\log_a(b/c)\]

OpenStudy (jiteshmeghwal9):

i have given the solution to ur first question already ;)

OpenStudy (anonymous):

thank you thank you

OpenStudy (jiteshmeghwal9):

Yw :)